OpenStudy (happyfeet8040):

Please help me?! Simplifysquare root of 3 multiplied by the fifth root of 3. 3 to the power of 1 over 10 3 to the power of 3 over 5 3 to the power of 9 over 10 3 to the power of 7 over 10

OpenStudy (happyfeet8040):

@dcabeche

OpenStudy (anonymous):

\[\sqrt{3} * \sqrt[5]{3} = (3) ^{1/2} * (3)^{1/5}\] \[(3)^{\frac {5}{10}}* (3) ^ {\frac {2}{10}} = (3) ^ {\frac {7}{10}}\]

OpenStudy (anonymous):

This is because \[a^n * a^m = a^{n+m}\]

OpenStudy (happyfeet8040):

Ugh, you're such a life saver. Thank you so much for all your help! :))

OpenStudy (anonymous):

No problem